OpenStudy (anonymous):

Emelina wrote the equation of a line in point-slope form as shown below. mc017-1.jpg What is Emelina’s equation in slope-intercept form? y = 3x + 2 y = 3x +10 y = 3x – 4 y = 3x – 10

OpenStudy (jack1):

y = mx + c m = gradient c = y-intercept y+4 = 3x+ 6 y= 3x +6 - 4 y = 3x +2